if the sum of a number and seven is tripled, the result is two less tha twice the number. find the number?
Answer by ankor@dixie-net.com(12701)

Just write an equation for what it says (x= "a number")
:
the sum of a number and seven is tripled, the result is two less the twice the number.
3(x+7) = 2x - 2
:
3x + 21 = 2x - 2
:
3x - 2x = -2 - 21
:
x = -23
;
;
Is this true? check in original equation:
3(-23+7) = 2(-23) - 2
3(-16) = -46 - 2
-48 = - 48; guess it is